# The 4DLO and other tubing models of $S^{3}$ symmetry

### v1

@@ Abstract
The Four-dimensional Light Orchestra or 4DLO was an interactive sculpture at the National Museum of Mathematics (MoMath) from November 20, 2025 through January 2026, illustrating various sub-symmetries of the 24-cell with colored lights, part of a larger sequence of tubing sculptures aiming to bring to life a few lines of tables appearing in [3], reprinted in [4], further illuminated in [11]. Best of all museum patrons could manipulate 4DLO’s lighting by singing and making funny noises into a microphone, and they did with gusto. Here we describe some of the technical aspects of this sculpture and its context.
Figure 1: (left) 4DLO showing a compound of three tesseracts, with student built work
just visible in the back. (right) Singing affected the sculpture and its program.
@@
Figure 2: (a) The fully symmetries of 24-cell; (b) half those, directing the edges; (c) A transitional moment during a sequence showing off the twenty four octahedral cells of the 24-cell, one by one. A similar sequence showed off the twenty four cubes of the three tesseracts sharing these edges. (d) Four mutually perpendicular groups of four rings of length six.
The 24-cell, also known as an “octaplex” and by other names, is a “four-dimensional regular polyhedron” or regular polytope, described more fully in a moment, denoted $\{3,4,3\}$ by Ludwig Schläfli who enumerated all regular polytopes in all dimensions in 1852. These are akin to the five regular (three-dimensional) polyhedra and to the infinitely many regular polygons in the plane. In each dimension $d\geq 5$ there are disappointingly only three regular polytopes, generalizations of the triangle, tetrahedron, etc., (denoted $\{3,...,3\}$ with one or more 3’s); the square, cube, tesseract, etc., ($\{...,3,4\}$ with no or more 3’s); and the polytopes dual to those, a tilted square, an octahedron, the 16-cell, etc. ($\{4,3,$...\} with no or more 3’s). In four dimensions, for whatever reasons, there are six.
Their enumeration was rediscovered repeatedly; George Hart has a helpful survey (with beautiful 3D printed models and a full bibliography) at [6], H.S.M. Coxeter’s Regular Polytopes [5] is a canonical reference, and the Wikipedia article is extensive with detailed technical information [15]. Schleimer and Segerman’s 2012 Bridges paper [13] gives precise definitions of many terms we use here.
Just as a polyhedron has two-dimensional polygons for faces, and can be considered as a division of the three dimensional sphere, a (four-d) polytope is formed of three-dimensional cells and may be viewed as a division of the hypersphere $S^{3}$. In the 24-cell there are twenty four three-dimensional octahedral cells. Each octahedron (denoted $\{3,4\}$) is composed of triangles $\{3\}$ meeting as though coming together through the sides of a square $\{4\}$. In a similar manner, at each vertex of the 24-cell, there are six octahedral cells, meeting tip to tip, as though they were coming through the faces of a cube $\{4,3\}$. Schläfli’s notation encodes this data as $\{3,4,3\}$.
@@
Clearly the 24-cell is an appealing shape. Mathematically it is highly symmetrical and an outlier among the regular polytopes, with no clear analogue in other dimensions. Like the tetrahedra-type polytopes of all dimensions, the 24-cell is self-dual — at the center of each of its twenty four cells there is a vertex of another 24-cell, its dual, and vice versa. For the same reason, it has the same number of two dimensional faces as it has edges, 96. Because all these polytopes have palindromic Schläfli symbols, and reversing the symbol of a regular polytope produces its dual, these are all duals of themselves, as are all of the regular polygons, which have symbols of the form $\{n\}$. Two copies of any of these self-dual polytopes form a natural compound. For example, the compound of two self-dual three-dimensional tetrahedra is well-known as Kepler’s Stella Octangula and two dual triangles form the familiar Star of David. So too is there a natural compound of two dual copies of the 24-cell — this structure appears in one of Rinus Roelef’s steel sculptures, landing a photo of the first author and Edmund Harriss in a newspaper [12]. Among these self-dual compounds, only the 24-cell has much structure to its symmetries.
@@
The symmetries of the hypersphere were enumerated in stages, begun by Goursat (1889), continued by Threlfall and Seifert (1931-3), Hurley (1951), Du Val (1964), and Conway and Smith (2003 [3]). In their recent monograph, Rastanawi and Rote give a geometric re-enumeration, which we lean on heavily for evaluating symmetry types [11]. (Their Section 3.3 gives a detailed survey of this history.) In effect, this project is to illustrate part of Conway and Smith’s tables decades late for The Symmetries of Things [4].
@@
Figure 3: (left) A spherical octahedron and its stereographic projection. Both are divisions of their space into triangular cells bounded by circular arcs meeting at right angles. (Right) A stereographic projection of a 16-cell, with tetrahedral cells and faces and edges meeting at right angles.
@@
The problem of visualizing the hypersphere is fundamentally three-dimensional, just as features on the surface of the Earth may be located with just two dimensions (say, latitude and longtitude). Nearly two thousand years ago Claudius Ptolemy produced the first atlas, with thousands of located sites spanning from the West African coast to the South China Sea. He projected points on Earth to a map, that is, from points on the sphere to points in a plane. His conical and cylindrical projections are foundational to cartographers to this day, but they’ve found little use for his stereographic projection despite it being very good for small regions away from some chosen point $N$. It is beloved by mathematicians, however, for not only does it preserves angles as measured on the sphere when they are measured on the plane, but it also takes circles on the sphere to circles in the plane (understanding, as did Nicolaus of Cusa ca. 1450, that lines are simply circles of infinite radius). And this mapping is elegantly defined: Geometrically, we choose some plane $P$ and a point $N$ on a sphere. Stereographic projection takes each point $p$ on the sphere to a point $\pi(p)$ in $P$, defined as the point where the line through $p$ and $N$ meets $P$. ($\pi(N)$ is itself “at infinity,” a kind of NaN.) In Figure 3 this is shown in the standard way, with $N$ at the top and $P$ tangent at N-N. In coordinates, say p=(x,y,,z,w)p=(x,y,...,z,w) of any dimension $d$, it is convenient to take N=(0,..,0,1)N=(0,..,0,-1) and project to the plane with $w=0$, using $\pi(p)=(x,...,z)/(1+w)$, a (d1)(d-1)-dimensional coordinate in $P$.
@@
By examining a pattern in $P$, such as the division of the plane into eight right-angled triangular regions at left in Figure 3, we can understand at least something about a division of the sphere in the same way (though a spherical octahedron will have much more symmetry than any planar drawing of it). We, as have many others, apply the same strategy one dimension up to learn something of the possible symmetries of patterns in the hypersphere.
Figure 4: (left and middle) Two cell-down 24-cells, shown as a compound of three tesseracts and as $T^{*}$ shown as three-fold rotational symmetries of a tetrahedron. [New pix coming] These models [will be] were built at the seventeenth Gathering for Gardner, Feb. 2026, with Scott Vorthmann.
@@
Tubing is manufactured in circular coils, and therefore pieces of it relax into circular arcs — this insight, that tubing is great for stereographic renderings of circles in the hypersphere, was the aha moment that spurred on this entire multifaceted project. With no delay the first author sought what might be at hand, finding some vinyl tubing at an automotive store and six-way connectors as playing jacks at a party store. The first of these models, a 16-cell, is at right in Figure 3, and the second is at left bottom in Figure 5.
@@
(Right) A Compound of Two Compounds of Three Sixteen Cells, G4G12 (2016).
We describe coordinates as quaternions, indicating $(a,b,c,d)$ as ${\bf{q}}=a{\bf{i}}+b{\bf{j}}+c{\bf{k}}+d{\bf{1}}$. On these we allow the usual rules of arithmetic, with multiplicative identity $\bf{1}$}, plus ${\bf{i}}{\bf{j}}={\bf{k}}$ and the rule that, for any imaginary (that is, ${\text{re}}({\bf{q}}):=d$ is zero) and unit ($a^{2}+b^{2}+c^{2}+d^{2}=1$) quaternion $\bf{q}$}, we have ${\bf{q}}{\bf{q}}=-{\bf{1}}$. (See [13] for further details.) The unit quaternions are exactly the points of the hypersphere $S^{3}$, and therefore coordinatize it. For illustrations, stereographically projecting from $\bf{1}$} into our space places $\bf{1}$} at the origin and $\bf{1}$} “at infinity”, the unit imaginary quaternions on the unit sphere and $\pm{\bf{i}},\pm{\bf{j}}$ and $\pm{\bf{k}}$ upon it as coordinate axes. An active reader might sketch on such a template. The familiar “cell-down” tesseract $\{4,3,3\}$ has sixteen vertices, we’ll call $V_{16}$, at points of the form $\frac{1}{2}(\pm{\bf{i}}\pm{\bf{j}}\pm{\bf{k}}\pm{\bf{1}})$, which in stereographic projection lie at the corners of two nested cubes to either side ($\pm{\bf{1}}$) of the unit sphere. We’ll further sort these points into the “even” ones $V_{16}^{+}$ with an even number of -’s, and its “odd” complement $V_{16}^{-}$. The edges of a tesseract connect vertices that differ by exactly one coordinate, and therefore each edge has one end in $V_{16}^{+}$ and the other in $V_{16}^{-}$. Three tesseracts can be arranged as at left in Figure 1; the yellow one, though tilted to the floor, is cell-down (a cube is at its center) and has vertices in $V_{16}$. It is dual to a “vertex-down” 16-cell with eight vertices, $V_{8}=\{\pm{\bf{1}},\pm{\bf{i}},\pm{\bf{j}},\pm{\bf{k}}\}$. You can quickly sketch a 16-cell of your own by connecting up these points on your template; or see the red 16-cell at far upper left Figure 5. A “cell-down” 16-cell with vertices in $V_{16}^{+}$ is at right in Figure 3 and is lit up at right in Figure 9. In fact $V_{16}\cup V_{8}$, which we denote $V_{24}$, form the vertices a 24-cell, which we’ll call “vertex-down”; its dual “cell-down” 24-cell has vertices $V_{24}^{\prime}$ of the form $\frac{1}{\sqrt{2}}(\pm{\bf{x}}\pm{\bf{y}})$ where ${\bf{x}},{\bf{y}}$ are two of ${\bf{1}},{\bf{i}},{\bf{j}}$, and $\bf{k}$}. The 4DLO of Figures 1 and 2 is vertex-down, but with the coordinate axes tilted in space. The 24-cells of Figure 4, [to be] made at G4G in February 2026, are cell-down.
@@
Figure 6: (left) an obliquely arrangement of three 16-cells; (middle) with directed edges, symmetry $O^{*}$; (right) a pleasing portion of three 16-cells made from four-way and two-way connectors.
[…]
@@
Nicely, quaternionic multiplication can describe symmetries like these: Each given unit quaternion ${\bf{q}}\in S^{3}$ also acts upon on $S^{3}$ as a whole, sliding everything together along a twisting family of great-circle rails (a “Hopf fibration”), just by multiplying every point by $\bf{q}$}. However multiplying on the left (${\bf{x}}\to{\bf{q}}{\bf{x}}$) slides along rails in right-handed [sic] coils and multiplying on the right (${\bf{x}}\to{\bf{x}}{\bf{q}}$) slides along left-handed coils (as measured against a right-handed ordering of $\bf{i}$}, $\bf{j}$}, and $\bf{k}$}). We can think of these coiling around two special “straight” rails, the one passing through $\bf{1}$} and $\bf{q}$}, and one opposite one, the furthest one away (these are special only in our coordinates though and geometrically all the opposite pairs of rails are equal). As $S^{3}$ slides one full circuit along a special rail, it makes a full revolution around it (the same thing as making a full circuit along its opposite rail).
@@
Or we may specify a fibration of $S^{3}$ into great circle rails by giving some unit $\bf{q}$} on the rail through $\bf{1}$}. We can choose that $\bf{q}$} to be imaginary, and therefore on the unit great sphere midway between $\bf{1}$} and $\bf{1}$}. Multiplying by ${\bf{r}}:=(\cos\theta){\bf{1}}+(\sin\theta){\bf{q}}$ moves the hypersphere $\frac{\theta}{2\pi}$th the way along this rail, but twisting with a handedness: Right multiplication ${\bf{x}}\to{\bf{x}}{\bf{r}}$ rotates $S^{3}$ by $\theta$ clockwise, as seen from $\bf{r}$} looking to $\bf{1}$}; and left-multiplication ${\bf{x}}\to{\bf{r}}{\bf{x}}$ rotates $S^{3}$ by $\theta$ counter-clockwise. Therefore taking ${\bf{x}}\to{{\bf{r}}}^{-1}{\bf{x}}{\bf{r}}$ slides $S^{3}$ along, then back, fixing $\bf{1}$} but rotating $S^{3}$ about the $\bf{1}$}-to-$\bf{r}$} axis clockwise by $2\theta=2\arccos({\text{re}}({\bf{r}}))$. We can thusly encode three-dimensional rotations by unit quaternions $\bf{r}$}. The correspondence is two-to-one, though, because ${{\bf{r}}}^{-1}{\bf{x}}{\bf{r}}={(-{\bf{r}})}^{-1}{\bf{x}}(-{\bf{r}})$ and so $\bf{r}$} and $\bf{r}$} give the same rotation of $\bf{1}$}. Each 3-d rotational symmetry group $G$ lifts to a symmetrical set points $G^{*}$ in $S^{3}$ (in binary, opposite pairs). Multiplying on the right (say) by any of these points $\bf{q}$} rotates $S^{3}$ taking $\bf{1}$} to $\bf{q}$} and preserving $G^{*}$ as a whole — $G^{*}$ is a fixed-point free symmetry group of $G^{*}$ itself. V. Hart et al. cleverly show these “binary” rotation groups of the platonic solids with monkeys [7].
As an example, orient a tetrahedron so that its vertices are $(1,1,1)$ and permutations of (1,1,1)(-1,-1,1), (that is, every other vertex of the unit cube) and let $T$ be its rotational symmetry group. It happens that the twenty-four points $V_{24}$ are the binary tetrahedral group $T^{*}$: The four pairs of $V_{8}$ correspond to the identity and the three 2-fold rotations, and the eight pairs of $V_{16}$ correspond to the eight 3-fold rotations. In fact, edges correspond to rotations: If ${\bf{p}},{\bf{r}}\in V_{24}$ are the ends of an edge then ${\bf{e}}:=({\bf{p}}{\bf{r}}^{-1})$ corresponds to a three-fold rotation (that is ${\bf{e}}^{3}=-{\bf{1}}$), and vice versa. Moreover rotations that have the same axis correspond to edges lying on the same great circle paths. Figure 2 (d) shows this on 4DLO. At right in Figure 4 we color cycles of 24-cell edges by which three-fold rotation axis they correspond to, showing symmetry $T^{*}$.
We can map the 24-cell to its dual, taking $V_{24}$ to $V_{24}^{\prime}$ by multiplying (right or left) by $\frac{1}{\sqrt{2}}({\bf{1}}+{\bf{i}})$, or by any other point in $V_{24}^{\prime}$. The twelve pairs of points in $V_{24}^{\prime}$ each correspond to a quarter turn about a coordinate axis, exactly the remaining rotational symmetries $O$ of a cube, and $O^{*}=V_{24}\cup V_{24}$. Connecting these vertices by edges corresponding to four-fold rotations we have the model at lower left in Figure 5, or if we color these by rotation axis, we have the one at right in Figure 7. Directing these edges gives symmetry $O^{*}$, but these models have more symmetry, though, because we can pivot around points:

### v2

@@ Abstract
The Four-dimensional Light Orchestra or 4DLO was an interactive sculpture at the National Museum of Mathematics (MoMath) from November 20, 2025 through January 2026, illustrating various sub-symmetries of the 24-cell with colored lights. This was part of a larger sequence of tubing sculptures aiming to bring to life a few lines of tables appearing in [3], reprinted in [4], and further illuminated in [11]. Best of all museum patrons could manipulate 4DLO’s lighting by singing and making funny noises into a microphone, and they did so with gusto. Here we describe some of the technical aspects of this sculpture and its context.
Figure 1: (left) 4DLO showing a compound of three tesseracts, with student built work just visible in back. (right) The colors and patterns of 4DLO responded to tone and rhythm.
@@
Figure 2: 4DLO showing (a) the full symmetries of 24-cell; (b) half those, directing the edges; (c) a sequence of the twenty four octahedral cells of the 24-cell, one by one (a similar sequence showed off the twenty four cubes of the three tesseracts sharing these edges); and (d) four mutually perpendicular groups of four rings of length six, indicated by color.
1  The 24-cell and other structures in the hypersphere, $S^{3}$
The 24-cell, also known as an “octaplex” and by other names, is a “four-dimensional regular polyhedron” or regular polytope, denoted $\{3,4,3\}$ by Ludwig Schläfli who enumerated all regular polytopes in all dimensions in 1852. These are akin to the five regular (three-dimensional) polyhedra and to the infinitely many regular polygons in the plane. In each dimension $d\geq 5$ there are disappointingly only three regular polytopes, generalizations of the triangle, tetrahedron, etc., (denoted $\{3,...,3\}$ with one or more 3’s); the square, cube, tesseract, etc., ($\{...,3,4\}$ with no or more 3’s); and the polytopes dual to those, a tilted square, an octahedron, the 16-cell, etc. ($\{4,3,$...\} with no or more 3’s). In four dimensions, as it turns out, there are six.
The list of regular polytopes was rediscovered many times. George Hart has a helpful survey (with beautiful 3D printed models and a full bibliography) at [6], H.S.M. Coxeter’s Regular Polytopes [5] is a canonical reference, and the Wikipedia article is extensive with detailed technical information. Schleimer and Segerman’s 2012 Bridges paper [13] gives precise definitions of many terms we use here.
Just as a polyhedron has two-dimensional polygons for faces, and can be considered as a division of the three-dimensional sphere, a four-dimensional polytope is formed of three-dimensional cells and may be viewed as a division of the hypersphere $S^{3}$. In the 24-cell there are twenty four three-dimensional octahedral cells. Each octahedron (denoted $\{3,4\}$) is composed of triangles $\{3\}$ meeting as though coming together through the sides of a square $\{4\}$. In a similar manner, at each vertex of the 24-cell, there are six octahedral cells, meeting tip to tip, as though they were coming through the square faces of a cube $\{4,3\}$. Schläfli’s notation encodes this data as $\{3,4,3\}$.
@@ 1  The 24-cell and other structures in the hypersphere, $S^{3}$
Clearly the 24-cell is an appealing shape. Mathematically it is highly symmetrical and an outlier among the regular polytopes, with no clear analogue in other dimensions. Like the tetrahedra-type polytopes of all dimensions, the 24-cell is “self-dual” — at the center of each of its twenty four cells there is a vertex of another 24-cell, its dual, and vice versa. For the same reason, it has the same number of two dimensional faces as it has edges, 96. All self-dual polytopes have palindromic Schläfli symbols because reversing the symbol of a regular polytope produces the symbol of its dual — regular polygons have symbols of the form $\{n\}$ and are also self-dual. Two copies of any of these self-dual polytopes form a natural compound. For example, the compound of two self-dual three-dimensional tetrahedra is well-known as Kepler’s Stella Octangula and two dual triangles form the familiar Star of David. So too is there a natural compound of two dual copies of the 24-cell — a portion of this structure appears in one of Rinus Roelef’s steel sculptures, landing a photo of the first author and Edmund Harriss in a newspaper [12]. Among these self-dual compounds, only the 24-cell has much structure to its symmetries (that is, structure to the subgroups of its symmetry group).
@@
The symmetries of the hypersphere were enumerated in stages, begun by Goursat (1889), continued by Threlfall and Seifert (1931-3), Hurley (1951), Du Val (1964), and Conway and Smith (2003, [3]). In their recent monograph [11], Rastanawi and Rote give a geometric re-enumeration, on which we lean heavily for evaluating symmetry types. (Their Section 3.3 gives a detailed survey of this history.) In effect, this project is to illustrate part of Conway and Smith’s tables decades late for The Symmetries of Things [4].
@@
Figure 3: At left, a spherical octahedron and its stereographic projection. Both are divisions of their space into triangular cells bounded by circular arcs meeting at right angles. Stereographic projections of 16-cells, with tetrahedral cells and faces and edges meeting at right angles: (middle) our first model made of tubing (right) a dynamic lit model in Fayetteville Ark, 2019. The polytopes in this figure are arranged “cell-down”, with a cell at the center of their projected image.
2  Rendering structures in the hypersphere
@@
The problem of visualizing the hypersphere is fundamentally three-dimensional, just as drawing pictures of features on the spherical Earth is inherently two-dimensional. Nearly two thousand years ago Claudius Ptolemy produced the first atlas, with thousands of located sites spanning from the West African coast to the South China Sea. He projected points on Earth to a map, that is, from points on the sphere to points in a plane. Cartographers find his conical and cylindrical projections foundational to this day, but they’ve found little use for his stereographic projection, despite it being very good for small regions away from some chosen point $N$. Mathematicians love it however, for not only does it preserve angles as measured on the sphere when they are measured on the plane, but it also takes circles on the sphere to circles in the plane (understanding, as did Nicolaus of Cusa ca. 1450, that lines are simply circles of infinite radius). And this mapping is elegantly defined: Geometrically, we choose some plane $P$ and a point $N$ on a sphere. Stereographic projection takes each point $p$ on the sphere to a point $\pi(p)$ in $P$, defined as the point where the line through $p$ and $N$ meets $P$. ($\pi(N)$ is itself “at infinity,” a kind of NaN.) In Figure 3 this is shown in the standard way, with $N$ at the top and $P$ tangent at N-N; we say “down” in this sense, towards this tangent. In coordinates, say p=(x,,z,w)p=(x,...,z,w) of any dimension $d$, it is convenient to take N=(0,..,0,1)N=(0,..,0,-1) and project to the plane with $w=0$, using $\pi(p)=(x,...,z)/(1+w)$, a (d1)(d-1)-dimensional coordinate in $P$.
@@
By examining a pattern in $P$, such as the division of the plane into eight right-angled triangular regions at left in Figure 3, we can understand at least something about a division of the sphere in the same way (though a spherical octahedron will have much more symmetry than any planar drawing of it). We, as have many others, apply the same strategy one dimension up to learn something of the possible symmetries in the hypersphere, like those of the “16-cells” that are shown middle and right in Figure 3.
@@
Tubing is manufactured in circular coils, and therefore pieces of it relax into circular arcs — this insight, that tubing is great for stereographic renderings of circles in the hypersphere, was the aha moment that spurred on this entire multifaceted project. With no delay the first author sought what might be at hand, finding some vinyl tubing at an automotive store and six-way connectors as playing jacks at a party store. The first of these models, a 16-cell, is at middle in Figure 3, and the second is at left in Figure 6.
Figure 4: (Left) With some deletions and distortions, stereographically projected sets $V_{8}$ (black), $V_{16}^{+}$ (blue), and $V_{16}^{-}$ (brown); with edges of a cell-down tesseract (black); the red points on the faces of this tesseract are the points in $V_{24}^{-}$. It can help to sketch your own version of this diagram. (Right) a screenshot of our interactive javascript model is at chaimgoodmanstrauss.com/4DLO.
Figure 5: Cell-down 24-cells, shown as (left) a compound of three tesseracts and (right) $T^{*}$ generated by the three-fold rotational symmetries of a tetrahedron. As with many of these models, not all edges of the mathematical structure are shown. The models in this figure are missing edges outside a unit ball, the “top” half of the 24-cell. Scott Vorthmann led a group build of these models at the sixteenth Gathering for Gardner, G4G16 (Feb. 2026).
@@
Figure 7: (Left) an obliquely arranged compound of three 16-cells. (Middle) if a direction were added to these edges, this model would show symmetry $O^{*}$. (Right) a pleasing piece of yard art, a portion of three 16-cells made from four-way and two-way connectors.
We describe coordinates as quaternions, indicating $(a,b,c,d)$ as ${\bf{q}}=a{\bf{i}}+b{\bf{j}}+c{\bf{k}}+d{\bf{1}}$. On these we allow the usual rules of arithmetic, with multiplicative identity $\bf{1}$}, plus the rules that ${\bf{i}}{\bf{j}}={\bf{k}}$ and, for any imaginary (that is, ${\text{Re}}({\bf{q}}):=d$ is zero) and unit ($a^{2}+b^{2}+c^{2}+d^{2}=1$) quaternion $\bf{q}$}, we have ${\bf{q}}{\bf{q}}=-{\bf{1}}$. (See [13] for further details.) The unit quaternions are exactly the points of the hypersphere $S^{3}$, and therefore coordinatize it. For illustrations, stereographically projecting from $\bf{1}$} into our space places $\bf{1}$} at the origin and $\bf{1}$} “at infinity”, the unit imaginary quaternions on the unit sphere and $\pm{\bf{i}},\pm{\bf{j}}$ and $\pm{\bf{k}}$ upon it as coordinate axes. Our sketch is at left in Figure 4 and the active reader is encouraged to draw their own. The familiar “cell-down” tesseract $\{4,3,3\}$ sits cell-downwards in $R^{4}$. It has sixteen vertices, we’ll call $V_{16}$, at points of the form $\frac{1}{2}(\pm{\bf{i}}\pm{\bf{j}}\pm{\bf{k}}\pm{\bf{1}})$. In stereographic projection these lie at the corners of two nested cubes to either side ($\pm{\bf{1}}$) of the unit sphere (and a cell is at the center of the projection). We’ll further sort these points into the “even” ones $V_{16}^{+}$ with an even number of -’s, and its “odd” complement $V_{16}^{-}$. The edges of a tesseract connect vertices that differ by exactly one coordinate, and therefore each edge has one end in $V_{16}^{+}$ and the other in $V_{16}^{-}$. Three tesseracts can be arranged as at left in Figure 1; the yellow one, though tilted to the floor, is cell-down (a cube is at its center) and has vertices in $V_{16}$. It is dual to a “vertex-down” 16-cell with eight vertices, $V_{8}=\{\pm{\bf{1}},\pm{\bf{i}},\pm{\bf{j}},\pm{\bf{k}}\}$ (One appears in red at the center of Figure 6; you can quickly sketch a vertex-down 16-cell by connecting up the points in $V_{8}$ on your drawing.) “Cell-down” 16-cells with vertices in $V_{16}^{+}$ are shown in Figure 3. In fact $V_{16}\cup V_{8}$, which we denote $V_{24}$, form the vertices a 24-cell, which we’ll call “vertex-down”; its dual “cell-down” 24-cell has vertices $V_{24}^{\prime}$ of the form $\frac{1}{\sqrt{2}}(\pm{\bf{x}}\pm{\bf{y}})$ where ${\bf{x}},{\bf{y}}$ are two of ${\bf{1}},{\bf{i}},{\bf{j}}$, and $\bf{k}$}. The 4DLO of Figures 1 and 2 show a vertex-down 24-cell (and thus with a vertex at the center of the projection), but with coordinates tilted so that gravity pulls in the (1,0,1)(-1,0,-1) direction. The 24-cells of Figure 5, made at the sixteenth Gathering For Gardner in February, 2026, are cell-down.
The vertex-down 24-cell shares its vertices $V_{24}$ with a compound of three 16-cells, with vertices $V_{8},V_{16}^{+}$ and $V_{16}^{-}$, and thus there is another compound of three 16-cells with corresponding vertices in $V_{24}^{\prime}$. We show these in Figures 6 and 7. In each compound of three 16-cells, the vertices of one 24-cell are those of the 16-cells, but the vertices of the dual 24-cell are merely points where the edges of each 16-cell appear to cross, just as edges of dual tetrahedra cross in a stella octangula. The two compounds of three 16-cells use the same edges but swap the role of the vertices! (The model at left in Figure 6 is both together overlapping.) Together these two compounds form A Compound of Two Compounds of Three 16-cells, assembled in two physical pieces in Atlanta in 2016 by participants of the twelfth biennial Gathering for Gardner, one of a sequence of sculptures we were fortunate to have created under the inspiring, supportive and generous leadership of George Hart, familiar to many of us in the math-art community.
There is a very nice compound of three tesseracts, dual to the compound of three 16-cells, sharing vertices $V_{24}$, which we show at left in Figures 1 and 5. (We don’t show the natural compound of six tesseracts.) You can find plans for making your own models in Figure 8.
Figure 8: Plans for a vertex-down projection of the 24-cell (brown); a cell-down 24-cell (light blue); and a compound of two compounds of three 16-cells (black); The points of $V_{24}$ are shaded and those of $V_{24}^{\prime}$ circled. These plans are on planes meeting the unit cube (of width 2) as shown.
3  Quaternionic groups as the symmetries of $S^{3}$
@@
Nicely, quaternionic multiplication can describe the symmetries of objects in $S^{3}$. Each given unit quaternion ${\bf{q}}\in S^{3}$ also acts upon on $S^{3}$ as a whole, sliding everything together along a twisting family of great-circle rails (a “Hopf fibration”), just by multiplying every point by $\bf{q}$}. However multiplying on the left (${\bf{x}}\mapsto{\bf{q}}{\bf{x}}$) slides along rails in right-handed [sic] coils and multiplying on the right (${\bf{x}}\mapsto{\bf{x}}{\bf{q}}$) slides along left-handed coils (as measured against a right-handed ordering of $\bf{i}$}, $\bf{j}$}, and $\bf{k}$}). We can think of these coiling around two special “straight” rails, the one passing through $\bf{1}$} and $\bf{q}$}, and its opposite, the one furthest one away and perpendicular to it. (These are special only in our coordinates: geometrically every rail has an opposite and any opposite pair of rails can be rotated to any other.) As $S^{3}$ slides one full circuit along a special rail, it makes a full revolution around it (the same thing as making a full circuit along its opposite rail).
Or we may specify a fibration of $S^{3}$ into great circle rails by giving some unit $\bf{q}$} on the rail through $\bf{1}$}. We can choose that $\bf{q}$} to be imaginary, and therefore on the unit great sphere midway between $\bf{1}$} and $\bf{1}$}. Multiplying by ${\bf{r}}:=(\cos\theta){\bf{1}}+(\sin\theta){\bf{q}}$ moves the hypersphere $\frac{\theta}{2\pi}$th the way along this rail, but twisting with a handedness: Right multiplication ${\bf{x}}\mapsto{\bf{x}}{\bf{r}}$ rotates $S^{3}$ by $\theta$ clockwise, as seen from $\bf{r}$} looking to $\bf{1}$}; and left-multiplication ${\bf{x}}\mapsto{\bf{r}}{\bf{x}}$ rotates $S^{3}$ by $\theta$ counter-clockwise. Therefore taking ${\bf{x}}\mapsto{{\bf{r}}}^{-1}{\bf{x}}{\bf{r}}$ slides $S^{3}$ along, then back, fixing $\bf{1}$} but rotating $S^{3}$ about the $\bf{1}$}-to-$\bf{r}$} axis clockwise by $2\theta=2\arccos({\text{Re}}({\bf{r}}))$. We can thusly encode three-dimensional rotations by unit quaternions $\bf{r}$}. The correspondence is two-to-one, though, because ${{\bf{r}}}^{-1}{\bf{x}}{\bf{r}}={(-{\bf{r}})}^{-1}{\bf{x}}(-{\bf{r}})$ and so $\bf{r}$} and $\bf{r}$} give the same rotation of $\bf{1}$}. Each 3D rotational symmetry group $G$ lifts to a symmetrical set of points $G^{*}$ in $S^{3}$ (in binary, opposite pairs). Multiplying on the right (say) by any of these points $\bf{q}$} rotates $S^{3}$ taking $\bf{1}$} to $\bf{q}$} and preserving $G^{*}$ as a whole — $G^{*}$ is a fixed-point free symmetry group of $G^{*}$ itself. V. Hart et al. cleverly show these “binary” rotation groups of the platonic solids using monkeys [7].
The 24-cell is related to the symmetries of the tetrahedron in this way (and a dual pair of 24-cells is related to the symmetries of a cube). To see this orient, a tetrahedron so that its vertices are $(1,1,1)$ and permutations of (1,1,1)(-1,-1,1) (that is, alternating vertices of the unit cube), and let $T$ be its rotational symmetry group. The four pairs of antipodal points in $V_{8}$ correspond to the identity and the three 2-fold rotations, and the eight pairs of antipodal points in $V_{16}$ correspond to the eight 3-fold rotations. In fact, edges correspond to rotations: If ${\bf{p}},{\bf{r}}\in V_{24}$ are the ends of an edge then ${\bf{e}}:=({\bf{p}}{\bf{r}}^{-1})$ corresponds to a three-fold rotation (that is ${\bf{e}}^{3}=-{\bf{1}}$), and vice versa. Moreover rotations that have the same axis correspond to edges lying on the same great circle paths. Figure 2(d) shows this on 4DLO. At right in Figure 5 we color cycles of 24-cell edges by which three-fold rotation axis they correspond to, showing symmetry $T^{*}$.
We can map the 24-cell to its dual, taking $V_{24}$ to $V_{24}^{\prime}$ by multiplying (right or left) by $\frac{1}{\sqrt{2}}({\bf{1}}+{\bf{i}})$, or by any other point in $V_{24}^{\prime}$. The twelve pairs of points in $V_{24}^{\prime}$ each correspond to a quarter turn about a coordinate axis, exactly the remaining rotational symmetries $O$ of a cube. The binary octahedral group $O^{*}=V_{24}\cup V_{24}^{\prime}$. Connecting these vertices by edges corresponding to four-fold rotations we have the model at left in Figure 6, or if we color these by rotation axis, we have the one at right in Figure 9. Directing these edges gives symmetry $O^{*}$, but these models have more symmetry, because we can pivot around points: By multiplying on both left and right, we can fix points, and therefore have symmetry about them.